# From Rows to Reasoning: A Retrieval-Augmented Multimodal Framework for Spreadsheet Understandingthanks: FRTR-Bench dataset: https://github.com/AnmolGulati6/FRTR-bench

### v1

@@ 5.1.1 Goal
The first experiment evaluates FRTR on the newly introduced FRTR-Bench dataset, assessing its ability to handle large, multimodal, and cross-sheet enterprise workbooks. We compare FRTR to SpreadsheetLLM, the strongest prior single-sheet compression baseline, across six large language models (LLMs) and three difficulty tiers (Easy, Medium, Hard). The objective is to measure reasoning accuracy, token efficiency, and latency in realistic spreadsheet settings that include multimodal content and multi-sheet dependencies.
5.1.2 Experimental Setup
@@
Both systems were tested on identical workbooks and queries drawn from FRTR-Bench. Each query was evaluated for correctness, reasoning trace, and provenance alignment. FRTR employed hybrid lexical–dense retrieval with RRF using the Amazon Titan Multimodal encoder, while SpreadsheetLLM followed its published compression–encoding approach [10]. All experiments were executed under identical prompt templates and model configurations.
5.1.3 Results Analysis
@@
Table 2 summarizes the comparative performance of FRTR and SpreadsheetLLM on FRTR-Bench. Across all difficulty levels, FRTR consistently outperforms SpreadsheetLLM. On easy workbooks, FRTR attains mean accuracies between 0.86 and 0.93 for top-tier models, exceeding SpreadsheetLLM’s 0.18–0.23 by more than two times. Performance remains robust on medium (0.62–0.65) and hard (0.66–0.72) tiers, where SpreadsheetLLM collapses below 0.10. Figure 2 displays the average results by difficulty across both FRTR and SpreadsheetLLM. Comprehensive results for each model by difficulty mode are reported in Appendix Appendix A. This confirms that compression-based methods do not scale when cross-sheet references and multimodal evidence are required, while the FRTR retrieval-based design preserves contextual fidelity at scale. FRTR also demonstrates high token efficiency: the average prompt length is only 7.8K tokens, compared to SpreadsheetLLM’s 13.1K tokens, corresponding to a mean compression ratio of 251× relative to raw workbook size. Despite smaller prompts, FRTR achieves higher accuracy, indicating that selective retrieval retains the most relevant evidence.
5.1.4 Discussion
The FRTR-Bench evaluation demonstrates that retrieval-first architectures outperform context-compression systems for large, multimodal spreadsheets. FRTR consistently achieves 2–4× higher accuracy 50–60% lower token usage, and comparable or lower latency than SpreadsheetLLM, while generalizing across diverse LLM families. Once the spreadsheet was indexed and embedded, retrieval time was negligible. These results validate FRTR-Bench as a challenging benchmark and establish FRTR as a state-of-the-art framework for advanced multimodal, retrieval-augmented spreadsheet understanding.
5.2 Experiment 2 - SpreadsheetLLM Benchmark Evaluation
5.2.1 Goal
@@ 5.2.2 Experimental Settings
We use the SpreadsheetLLM benchmark [10], which consists of single-sheet, text-only question–answer pairs requiring numerical reasoning, formula synthesis, and contextual lookup comprehension. Unlike FRTR-Bench, this dataset excludes multimodal and cross-sheet reasoning, emphasizing tabular text understanding within context-limited settings.
@@
Table 3: Performance comparison on the SpreadsheetLLM benchmark. FRTR achieves answer accuracy comparable to SpreadsheetLLM.
Technique Model Answer Accuracy Mean Tokens Latency (s)
FRTR GPT-4o 0.77 6,920 2.48
GPT-5 0.87 6,920 18.06
Claude 4.5 Sonnet 0.84 6,920 10.40
Gemini 2.5 Pro 0.85 6,920 24.06
LLaMA Maverick-17B-instruct 0.53 6,920 2.18
GPT-OSS-120b 0.75 6,920 7.37
SpreadsheetLLM GPT-4o 0.78 5,811 0.81
GPT-5 0.90 5,811 5.50
Claude 4.5 Sonnet 0.91 5,811 5.49
Gemini 2.5 Pro 0.89 5,811 10.00
LLaMA Maverick-17B-instruct 0.73 5,811 0.73
GPT-OSS-120b 0.80 5,811 4.01
Naïve Full-Context GPT-4o 0.35 13,631 3.01
GPT-5 0.68 13,631 18.07
Claude 4.5 Sonnet 0.55 13,631 9.06
Gemini 2.5 Pro 0.67 13,631 15.48
LLaMA Maverick-17B-instruct 0.41 13,631 2.45
GPT-OSS-120b 0.59 13,631 8.61
Baselines.
  1. 1.
  2. 2.
5.2.3 Results
Table 3 summarizes model performance across all methods.
5.2.4 Results Analysis
@@
As shown in Table 3, FRTR performs on par with SpreadsheetLLM in overall accuracy across most model families while substantially outperforming the naïve full-context baseline. For high-capacity models such as GPT-5, Claude 4.5 Sonnet, and Gemini 2.5 Pro, FRTR attains accuracies in the 0.84–0.87 range, closely matching SpreadsheetLLM’s 0.89–0.91. Open-weight models follow the same trend, achieving strong relative performance despite having smaller token counts. Importantly, FRTR achieves this parity while using roughly 40–50% fewer tokens than the full-context baseline and maintaining stable latency across models. The retrieval-based design also provides explicit provenance for every reasoning step, enabling interpretability that compression-based systems lack.
5.2.5 Discussion
@@
Although FRTR does not surpass SpreadsheetLLM in raw accuracy on this single-sheet benchmark, it achieves near-equivalent results while offering stronger efficiency and transparency benefits. It is important to further note that FRTR’s approach considered every sheet in an Excel workbook, while SpreadsheetLLM only compressed and provided the relevant sheet to the LLM. FRTR’s advanced multimodal retrieval-augmented approach focuses on retrieving the minimal evidence needed for reasoning rather than compressing the entire sheet, reducing token consumption and supporting explainable, auditable outputs. These findings indicate that FRTR generalizes effectively even in non-multimodal, single-sheet scenarios where its architectural advantages (cross-sheet retrieval and image reasoning) are not fully utilized. Matching SpreadsheetLLM’s performance under these constrained conditions highlights the robustness of retrieval-augmented reasoning and its potential as a scalable, interpretable alternative to context-compression approaches.
6 Conclusion
@@
Large Language Models (LLMs) continue to struggle with reasoning over large-scale, enterprise-grade spreadsheets that contain thousands of rows, multiple linked sheets, and multimodal elements such as charts and receipts. Existing approaches based on single-sheet compression or full-context encoding fail to scale and do not reflect how real users interact with complex analytical workbooks. To address these challenges, we introduced From Rows to Reasoning (FRTR), an advanced, multimodal retrieval-augmented generation framework that decomposes spreadsheets into granular row, column, and block embeddings, employs hybrid lexical–dense retrieval with Reciprocal Rank Fusion (RRF), and integrates multimodal embeddings to reason jointly over numerical and visual content. Alongside this framework, FRTR-Bench provides the first large-scale benchmark for multimodal spreadsheet reasoning, comprising 30 enterprise-grade Excel workbooks spanning nearly four million cells and more than 50 embedded images. We evaluated FRTR using six LLMs, achieving 74% answer accuracy on FRTR-Bench, exceeding prior state-of-the-art approaches by 50%. On the SpreadsheetLLM benchmark, FRTR achieved 87% while reducing token usage by roughly 50% compared to context-compression methods. These results show that retrieval-first architectures enable scalable, interpretable, and efficient spreadsheet reasoning. FRTR gives LLMs the critical ability to reason over massive spreadsheets in domains such as finance, auditing, and healthcare, where verifiability and efficiency are critical. Future work should explore adaptive retrieval, finer multimodal alignment, and agentic spreadsheet manipulation.
Appendix Appendix A Experiment 1 - FRTR-Bench Evaluation
Table 4: Detailed per-model results on FRTR-Bench grouped by difficulty tier. Times are in seconds; tokens are prompt tokens.
Difficulty Model FRTR SpreadsheetLLM Accuracy Tokens Latency (s) Accuracy Tokens Latency (s) Easy Gemini 2.5 Pro 0.864 12,088 21.38 0.205 6,673 20.97 GPT-5 0.864 12,088 13.29 0.227 6,673 14.60 Sonnet 4.5 0.932 12,088 10.79 0.182 6,673 7.12 GPT-4o 0.545 12,088 3.97 0.091 6,673 0.89 LLaMA Maverick-17B 0.750 12,088 2.25 0.205 6,673 0.90 GPT-OSS-120B 0.545 12,088 5.00 0.000 6,673 7.68 Medium Gemini 2.5 Pro 0.557 7,235 30.84 0.278 17,483 38.85 GPT-5 0.646 7,235 16.10 0.114 17,483 20.68 Sonnet 4.5 0.658 7,235 12.31 0.089 17,483 9.68 GPT-4o 0.443 7,235 5.08 0.101 17,483 1.33 LLaMA Maverick-17B 0.443 7,235 2.44 0.245 17,483 1.46 GPT-OSS-120B 0.468 7,235 6.59 0.148 17,483 9.93 Hard Gemini 2.5 Pro 0.694 3,318 24.99 0.194 9,173 35.98 GPT-5 0.722 3,318 16.88 0.250 9,173 16.70 Sonnet 4.5 0.667 3,318 11.53 0.167 9,173 8.75 GPT-4o 0.528 3,318 5.17 0.083 9,173 1.02 LLaMA Maverick-17B 0.583 3,318 2.25 0.222 9,173 0.78 GPT-OSS-120B 0.556 3,318 4.82 0.083 9,173 7.69
Appendix Appendix B Token Usage Visualization
Figure 3: Visualization of prompt token usage for FRTR and SpreadsheetLLM.
Appendix Appendix C Prompt Template: explain_prompt

### v2

@@ 5.1.1. Goal
The first experiment evaluates FRTR on the newly introduced FRTR-Bench dataset, assessing its ability to handle large, multimodal, and cross-sheet enterprise workbooks. We compare FRTR to SpreadsheetLLM, the strongest prior single-sheet compression baseline, across six large language models (LLMs) and three difficulty tiers (Easy, Medium, Hard). The objective is to measure reasoning accuracy, token efficiency, and latency in realistic spreadsheet settings that include multimodal content and multi-sheet dependencies. Since SpreadsheetLLM was designed for single-sheet reasoning, we extended it to multi-sheet settings by compressing each sheet individually and concatenating all compressed sheets into a single prompt. This concatenation fit within model context limits without truncation, representing the most favorable adaptation of their approach to cross-sheet tasks.
5.1.2. Experimental Setup
@@
Both systems were tested on identical workbooks and queries drawn from FRTR-Bench. Each query was evaluated for correctness, reasoning trace, and provenance alignment. FRTR employed hybrid lexical–dense retrieval with RRF using the Amazon Titan Multimodal encoder, while SpreadsheetLLM followed its published compression–encoding approach (Dong et al., 2025b). All experiments were executed under identical prompt templates and model configurations.
5.1.3. Results Analysis
@@
Table 2 summarizes the comparative performance of FRTR and SpreadsheetLLM on FRTR-Bench. Across all difficulty levels, FRTR consistently outperforms SpreadsheetLLM. On easy workbooks, FRTR attains mean accuracies between 0.86 and 0.93 for top-tier models, exceeding SpreadsheetLLM’s 0.18–0.23 by more than two times. Performance remains robust on medium (0.62–0.65) and hard (0.66–0.72) tiers, where SpreadsheetLLM collapses below 0.10. Figure 3 displays the average results by difficulty across both FRTR and SpreadsheetLLM. Comprehensive results for each model by difficulty mode are reported in Appendix A. This confirms that compression-based methods do not scale when cross-sheet references and multimodal evidence are required, while the FRTR retrieval-based design preserves contextual fidelity at scale. FRTR also demonstrates high token efficiency: the average prompt length is only 7.7K tokens, compared to SpreadsheetLLM’s 13.1K tokens, corresponding to a mean compression ratio of 251× relative to raw workbook size. Despite smaller prompts, FRTR achieves higher accuracy, indicating that selective retrieval retains the most relevant evidence.
5.1.4. Discussion
The FRTR-Bench evaluation demonstrates that retrieval-first architectures outperform context-compression systems for large, multimodal spreadsheets. FRTR consistently achieves 2–4× higher accuracy, 40% lower token usage, and comparable or lower latency than SpreadsheetLLM, while generalizing across diverse LLM families. Figure 3 demonstrates FRTR’s high accuracy across various spreadsheet lengths. The U-shaped accuracy pattern suggests that structural complexity, not size alone, drives difficulty. Very large spreadsheets in the benchmark are primarily transaction logs with repetitive structure and simple aggregation queries (e.g., payroll records, time-series data), enabling effective retrieval and reasoning. In contrast, medium spreadsheets often represent complex business scenarios (consolidations, multi-entity reporting) with multiple interconnected worksheets, cross-references, and semantically similar sections that create retrieval ambiguity. This ’confusion zone’ challenges both approaches: FRTR’s retrieval must disambiguate between multiple similar contexts, while SpreadsheetLLM’s compression discards critical cross-sheet relationships.
Notably, FRTR achieves 65% accuracy on the 23 image-dependent questions in FRTR-Bench, enabling a class of multimodal queries that text-only approaches cannot address. Once the spreadsheet was indexed and embedded, retrieval time was negligible and inference time (as reported in Table 2) was relatively consistent across approaches. These results validate FRTR-Bench as a challenging benchmark and establish FRTR as a state-of-the-art framework for advanced multimodal, retrieval-augmented spreadsheet understanding.
5.2. Experiment 2 - SpreadsheetLLM Benchmark Evaluation
5.2.1. Goal
@@ 5.2.2. Experimental Settings
We use the SpreadsheetLLM benchmark (Dong et al., 2025b), which consists of single-sheet, text-only question–answer pairs requiring numerical reasoning, formula synthesis, and contextual lookup comprehension. Unlike FRTR-Bench, this dataset excludes multimodal and cross-sheet reasoning, emphasizing tabular text understanding within context-limited settings. For SpreadsheetLLM’s evaluation on this benchmark, the compressed oracle sheet is provided as input, following the methodology detailed in SpreadsheetLLM (Dong et al., 2025b).
@@
Table 3. Performance comparison on the SpreadsheetLLM benchmark. FRTR achieves answer accuracy comparable to SpreadsheetLLM. Mean tokens: FRTR 6,920; SpreadsheetLLM 5,811; Naïve 13,631.
Technique Model Answer Accuracy Latency (s)
FRTR GPT-4o 0.77 2.48
GPT-5 0.87 18.06
Claude Sonnet 4.5 0.84 10.40
Gemini 2.5 Pro 0.85 24.06
LLaMA Maverick-17B-instruct 0.53 2.18
GPT-OSS-120b 0.75 7.37
SpreadsheetLLM GPT-4o 0.78 0.81
GPT-5 0.90 5.50
Claude Sonnet 4.5 0.91 5.49
Gemini 2.5 Pro 0.89 10.00
LLaMA Maverick-17B-instruct 0.73 0.73
GPT-OSS-120b 0.80 4.01
Naïve Full-Context GPT-4o 0.35 3.01
GPT-5 0.68 18.07
Claude Sonnet 4.5 0.55 9.06
Gemini 2.5 Pro 0.67 15.48
LLaMA Maverick-17B-instruct 0.41 2.45
GPT-OSS-120b 0.59 8.61
Figure 4. Unit granularity ablation: accuracy vs. latency, bubble size = token count (Claude Sonnet 4.5, FRTR-Bench). All Types (FRTR) achieves 0.74 accuracy with 7,690 tokens, outperforming SpreadsheetLLM with Gemini 2.5 Pro (0.24, 12,745 tokens).
5.2.3. Results Analysis
@@
As shown in Table 3, FRTR performs on par with SpreadsheetLLM in overall accuracy across most model families while substantially outperforming the naïve full-context baseline. For high-capacity models such as GPT-5, Claude Sonnet 4.5, and Gemini 2.5 Pro, FRTR attains accuracies in the 0.84–0.87 range, closely matching SpreadsheetLLM’s 0.89–0.91. Open-weight models follow the same trend, achieving strong relative performance despite having smaller token counts. Importantly, FRTR achieves this parity while using roughly 40–50% fewer tokens than the full-context baseline and maintaining stable latency across models. The retrieval-based design also provides explicit provenance for every reasoning step, enabling interpretability that compression-based systems lack.
5.2.4. Discussion
@@
Although FRTR does not surpass SpreadsheetLLM in raw accuracy on this single-sheet benchmark, it achieves near-equivalent results while offering stronger efficiency and transparency benefits. It is important to further note that FRTR’s approach considered every sheet in an Excel workbook, while SpreadsheetLLM was given the oracle sheet (as was implemented in the SpreadsheetLLM paper). FRTR’s advanced multimodal retrieval-augmented approach focuses on retrieving the minimal evidence needed for reasoning rather than compressing the entire sheet, reducing token consumption and supporting explainable, auditable outputs. These findings indicate that FRTR generalizes effectively even in non-multimodal, single-sheet scenarios where its architectural advantages (cross-sheet retrieval and image reasoning) are not fully utilized. Matching SpreadsheetLLM’s performance under these constrained conditions highlights the robustness of retrieval-augmented reasoning and its potential as a scalable, interpretable alternative to context-compression approaches.
5.3. Ablation Studies
To validate the design choices in FRTR, we conduct ablation experiments on FRTR-Bench using the highest performing LLM (Claude Sonnet 4.5), examining two key dimensions: retrieval methodology and unit granularity. In both cases, we retrieve the top 10 most relevant documents.
5.3.1. Retrieval Method Ablation
We compare three retrieval strategies on FRTR-Bench to assess the contribution of hybrid search: (1) Hybrid RRF combining lexical and dense retrieval, (2) Vector-only dense retrieval, and (3) BM25 lexical-only retrieval.
Table 4. Retrieval method ablation on FRTR-Bench. Results show answer accuracy, retrieval time, and tokens retrieved across different search strategies.
Retrieval Method Answer Accuracy Avg Retrieval Time (s) Avg Tokens
Hybrid (RRF) 0.74 0.82 6,903
Vector Only 0.59 0.70 5,857
BM25/Lexical Only 0.42 0.11 5,616
The results demonstrate that hybrid RRF retrieval achieves significantly higher accuracy (0.74) compared to vector-only (0.59) and lexical-only (0.42) approaches. While BM25 offers the fastest retrieval time (0.11s), its accuracy is insufficient for complex spreadsheet reasoning. Vector-only retrieval improves upon lexical search but still falls short of hybrid performance, suggesting that semantic similarity alone misses exact-match patterns captured by BM25. The hybrid approach combines the strengths of both methods through rank fusion, achieving the best accuracy with only a modest increase in retrieval time (0.82s). This validates FRTR’s design choice to employ RRF-based hybrid retrieval for robust, scalable spreadsheet understanding.
5.3.2. Unit Granularity Ablation
We evaluate different indexing granularities by comparing retrieval using only rows, columns, or windows against our default configuration, which combines all unit types. In all cases, images are included in the retrieval set. This ablation isolates the contribution of each granularity level to overall reasoning performance.
Table 5. Unit granularity ablation on FRTR-Bench. Results demonstrate the complementary value of multi-granular indexing for spreadsheet reasoning.
Units Retrieved Answer Accuracy Avg LLM Time (s) Avg Tokens
Rows 0.52 7.55 2,017
Columns 0.18 5.34 11,465
Windows 0.58 7.15 7,259
All Types 0.74 11.71 7,690
The results reveal that different granularities capture complementary information, as visualized in Figure 4. Row-only retrieval achieves 0.52 accuracy, insufficient for cross-column reasoning. Column-only retrieval performs poorly (0.18 accuracy), as column serialization lacks sufficient context and produces verbose representations. Window-based retrieval achieves 0.58 accuracy, capturing spatial relationships within localized regions.
Combining all unit types (rows + columns + windows + images) achieves the highest accuracy (0.74), demonstrating that multi-granular indexing surfaces relevant evidence regardless of query structure. Notably, FRTR’s multi-granular approach substantially outperforms the SpreadsheetLLM baseline (0.24 accuracy), achieving over 3× improvement while maintaining comparable token efficiency. While this incurs slightly higher latency (11.71s) due to increased retrieval and reasoning complexity, the accuracy gain justifies the cost. Token cost variation is due to benchmark characteristics: most spreadsheets had long tables with few columns, resulting in longer column documents and shorter row documents. This confirms that FRTR’s multi-granular design is essential for robust spreadsheet reasoning across diverse query types.
5.4. Cost Analysis
LLM inference costs scale with token consumption. Table 2 shows FRTR uses 7,691 tokens on average vs. SpreadsheetLLM’s 12,745 tokens, a 40% reduction. At typical API rates (\$0.01–0.03 per 1K input tokens for frontier models), this translates to significant savings at scale. For an enterprise processing 1,000 queries daily, FRTR would reduce annual LLM costs by approximately 40%, while improving accuracy. Embedding costs are incurred once during indexing and amortized across all subsequent queries.
6. Limitations
While FRTR demonstrates strong performance, certain design choices merit discussion. First, our multimodal embedding approach achieves 65% accuracy on image-dependent questions, a capability that would otherwise be zero without multimodal retrieval, and we anticipate further gains as enterprise-grade multimodal encoders continue to mature. Second, FRTR-Bench focuses on English-language spreadsheets, though the retrieval architecture is language-agnostic and extensible to multilingual settings. Third, FRTR generates formulas and cell references rather than executing them directly, preserving human oversight aligned with enterprise audit requirements. Finally, while retrieval latency is negligible once indexed (under 1 second), initial workbook indexing is a one-time cost amortized across all subsequent queries. The initially substantial indexing time was alleviated by employing batch processing and parallelization techniques.
7. Ethics and Reproducibility
FRTR-Bench consists of synthetic enterprise-style workbooks designed to simulate realistic analytical scenarios without using real user data or proprietary information. The benchmark is available via a github repository111 FRTR-Bench Repository (https://github.com/AnmolGulati6/FRTR-bench).
8. Conclusion
@@
Large Language Models (LLMs) continue to struggle with reasoning over large-scale, enterprise-grade spreadsheets that contain thousands of rows, multiple linked sheets, and multimodal elements such as charts and receipts. Existing approaches based on single-sheet compression or full-context encoding fail to scale and do not reflect how real users interact with complex analytical workbooks. To address these challenges, we introduced From Rows to Reasoning (FRTR), an advanced, multimodal retrieval-augmented generation framework that decomposes spreadsheets into granular row, column, and block embeddings, employs hybrid lexical–dense retrieval with Reciprocal Rank Fusion (RRF), and integrates multimodal embeddings to reason jointly over numerical and visual content. Alongside this framework, FRTR-Bench provides the first large-scale benchmark for multimodal spreadsheet reasoning, comprising 30 enterprise-grade Excel workbooks spanning nearly four million cells and more than 50 embedded images. We evaluated FRTR using six LLMs, achieving 74% answer accuracy on FRTR-Bench, achieving 50 percentage points higher accuracy than prior state-of-the-art approaches (74% vs. 24%). On the SpreadsheetLLM benchmark, FRTR achieved 87% accuracy while reducing token usage by roughly 50% compared to a full-context baseline. These results show that retrieval-first architectures enable scalable, interpretable, and efficient spreadsheet reasoning. FRTR gives LLMs the critical ability to reason over massive spreadsheets in domains such as finance, auditing, and healthcare, where verifiability and efficiency are critical. Future work should explore adaptive retrieval, finer multimodal alignment, and agentic spreadsheet manipulation.
Appendix A Experiment 1 - FRTR-Bench Evaluation
Appendix B Token Usage Visualization
Refer to caption Bar chart comparing prompt token usage between FRTR and SpreadsheetLLM across different spreadsheet files. FRTR uses significantly fewer tokens than SpreadsheetLLM due to its retrieval-based approach.
Figure 5. Visualization of prompt token usage for FRTR and SpreadsheetLLM.
Appendix C Prompt Template